Aggregator

APPROVED B-36 Replace damaged breakers

4 days 4 hours ago
Log Number: 4551312Logbook(s): Body: 

Via e-mail touched based with DC Power, MSS, and AES Electricians who indicated they'll be working in the Bldg today. RF comment below. All have confirmed this outage is compatible with their use of the building today.

Once the PTW is on-issue this task is approved for execution today, 10/29. Added to SAM Cal and Work Map.

moser

C100-02R Acceptance Testing Swing Shift

4 days 14 hours ago
Log Number: 4551309Logbook(s): Body: 

Completed Tuner characterization for cavity 4.
Then did the gradient ramp for cavity 4. Emax at 23.1 MV/m - quench limited.
FE Onset at ~18 MV/m
Max dose rate ~ 500 mR/hr at 22.6 MV/m
Attempted one hour at 22.6 MV/m. After about 45 minutes, klystrons faulted on KMAV.
Completed FE vs E measurement

drury

COMMENT ON 2026 SAM LCW Controls Calibrations

4 days 16 hours ago
Log Number: 4551307Logbook(s): Body: 

Calibration of North LINAC, West ARC, South LINAC, and East ARC dissolved oxygen sensors completed. Plan is to replace failed two North Access Honeywell UDC Thursday 10/29/26 and then move to the South Access to begin replacement of several failed Honeywell UDC's. Replacement UDCs arrived recently.

jonesc

MOLLER install 10/28/2026

4 days 18 hours ago
Log Number: 4551305Logbook(s): Tags: Entry Makers: jonesBody: 

• Leak checking the hydrogen piping.
• Prepared a cable tray support stand for welding.
• Upstream pivot platform girder support is aligned and ready for S&A.

jones

iocsoftrfcrydev1 restored

4 days 18 hours ago
Log Number: 4551303Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2026-10-28 16:29:08 The output from the myRestore execution was as follows: iocsoftrfcrydev1 systemEnum@2026-10-28 16:29 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -i iocsoftrfcrydev1 (success) PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 Procedure return status: Success
alarms

Follow-up Re: Cryo Daily Update

4 days 18 hours ago
Log Number: 4551302Logbook(s): Entry Makers: creelBody: 

Performed SCN CC4 spin test up to 70HZ
Oscilloscope traces of position sensors looked good.
No wobble identified. Position feedback signal amplitude did not grow with speed and bearing currents remained low.
More testing in the AM.

creel

Hall C daily summary

4 days 19 hours ago
Log Number: 4551299Logbook(s): Body: 

Today Hall C technicians worked on.

Getting Fall protection permit, Rad work permit and EPAS approved and signed.
Cross training on Bridgeport milling machine
EPAS training
ENGE mobile detector stand.
NUCAT scintillator frame holder.
Assembling filament extrusion project parts.

Thursday 10/29/26 we have requested RADCON support to remove down stream beam pipe.

jnines

iocsoftrfcrydev1 restored

4 days 19 hours ago
Log Number: 4551296Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2026-10-28 16:03:57 The output from the myRestore execution was as follows: iocsoftrfcrydev1 systemEnum@2026-10-28 16:03 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -i iocsoftrfcrydev1 (success) PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 Procedure return status: Success
alarms

prep chamber rework

4 days 19 hours ago
Log Number: 4551295Logbook(s): Entry Makers: marcyBody: 

Vacuum recovering - Overboard ion pump came on and is at 540 uA. We'll leave this on and try to turn on the chamber ion pump tomorrow for checkout of leak valve, Cs and RGA as well as tests of long manipulator.

Cleanroom still in place in case we have to rework any of the components.

marcy

Cryogenics Update

4 days 19 hours ago
Log Number: 4551292Logbook(s): Entry Makers: creelBody: 

Completed installation of updated CHL2 horizontal and vertical ASME relief valves.
SCN CC4 unlocked and levitated
SCN CC4 time response damping, air-gap, and UPS battery testing completed on magnetic bearing control cabinet
Continued removing unused racks and other hardware from CHL control room

creel

COMMENT ON Install 1L06,07 utubes

4 days 19 hours ago
Log Number: 4551291Logbook(s): Body: 

1L06 is online and cold. 1L07 was punted last week due to excessive pressure buildup when attempting to install the return U-Tube.
This morning, a purge rig was installed on 1L07, and we have been flowing through the module since.

johnw

1I beamline NEG activations

4 days 19 hours ago
Log Number: 4551288Logbook(s): Entry Makers: marcyBody: 

I activated the NEG pumps in the 1I beamline. This was last done when they were installed in 2021, so it was time for a reactivation anyway, and they saw a tiny leak due to the gun bakeout and a leak in the gun valve bonnet seal.

Since the pumps are distributed along ~10' of beamline, and I had a pump cart installed only in the upstream location, I activated them in a staggered start/staggered end, starting with the end farthest from the gun.

For reactivation, the NEG should be taken to about 75% of the temperature for full activation.

marcy

Shift Information

Shift Day
CEBAF Program
MCC Crew Chief Deir
MCC Operators
Program Deputy Delk
Announcements

Access Status

Recent PSS History

recent pss history graph